Fermi National Laboratory
 

Accelerator Update for September 7, 2007 - September 14, 2007

 

Fermilab Shutdown Report
9/7/07 – 9/14/07

LINAC

Things Done This Week

  • Completed This Week
    • Started installation of next LLRF VXI chassis on LRF3
    • Repaired Modulator Voltage Readback issue on LRF3
    • Installed Linac Filament Inductor motor and controller for use in manual control mode on LRF4
    • Continued NTF treatment on schedule
    • Will switch from I- to H- today
  • To do next week
    • NTF Treatment scheduled for Monday, Sept. 17 th
    • Finish VXI Installation on LRF3
    • Test new Linac Filament Inductor motor and controller in auto control regulation on LRF4
    • Clean dome and refurbish source on I-
    • Install new vacuum pump on I- Source
    • Fix H.E. Linac Debuncher Water skid issue with temperature regulation
  • 10 day Linac shutdown starts on Tuesday, Sept. 18th
    • After success of 9" line cavity coupler loop tuning on LRF3, we will repeat the procedure on LRF5 to reduce reverse during beam time. This will start on Tuesday, require beam pulses to verify beam loading / reverse power that night if all goes well, then return shielding blocks on Wednesday.
    • Start and Finish LLRF VXI Installation on LRF1 & LRF2
    • Install Co-axial cables needed for LLRF Controls
    • Install controls electronics in L.E. Linac

Booster

Water work:

  • LCW water work in the Booster West Gallery
    • Final water connections from the header to the various systems are complete
    • BRF17 and BRF18 manifold re-work will finish today
      • A final walk thru inspection with experts and interested parties will be arranged today before filling the system
    • Next week: Fill up the system and perform hydrostatic tests
  • CUB
    • Booster de-oxy skid is finished and the 95LCW water has been flowing through the system.
    • 95LCW pump upgrade is complete and operational.
  • Corrector installation
    • Number of new correctors in the tunnel: 11
      • L19, L17, L21, L23, L15, L9, L11, L13, L5, L7, & L1
      • A comprehensive plan has been worked out and adopted for the corrector installation at L3
        • One of the most challenging location for installation
        • Requires disassemble of some 400MeV components
        • Requires replacement of a dog leg magnet
        • Requires further shielding surrounding the area
        • Requires elaborate mechanical design
      • Dry run transporting the dogleg through tight clearances
      • Dismantling 400MeV components begin today!
    • Next week
      • Start L3 installation
  • Protective screen added to the corrector power supplies
  • Next week: Energize the correctors at low current, one at a time, for field measurements
  • Additional Activities
    • BET sinkhole has been repaired
    • Water leak developed in the BWT parking lot
      • FESS determined the leak comes from a ICW line from CUB
    • RF maintenance progressing well
    • Q800 replacement will proceed next week

Antiproton Source

Mechanical

  • Leak Check (4)
    • AP3 to be done
    • Finished checking D/A, A40, A30
      • D:ESEP, A30 IP fixed
  • D:ISEP Shielding
  • Ion Chamber removal
  • Deb DCCT motorized stand
    • Ready for alignment
  • D:EKIK ceramic tube
    • New tube on stand; ready for alignment
  • AP2 Water Hose Replacement
  • Leaking LQA
    • Surveyed into place
  • Replaced leaking Deb Sextuple
  • SMA water leak
    • Surveyed into place
  • 4 water leaks to fix
    • IQ8 fixed twice
  • A10 Obstruction Search
    • Opened; components to shuffle

Target Station

  • Water Systems (2)
  • Temperature Monitoring (2)
    • 1 wiring done
  • Target Rotation
  • Storage
  • SEM Module
  • LVDTs
  • Sample Retrieval
  • Lens motion
  • Air Exhaust Stack
  • 10mm-2 into transformer
    • Prepping; await tools
  • 10mm-3 into transformer
    • Prepping, installing
  • Lens replacement

Cryo

  • Valve Stem replace
  • Leak Check (1 of 2)
  • LN2 transfer line maintenance

Cooling Systems

  • Deb 3&4 Trans Notch Filters
  • Deb Double Notch Filters
  • Core 4-8GHz Equalizer
  • TWT balance & trip points
    • Cryo cold; testing 25% complete
  • Stacktail Filter 3 cable change
    • Cryo cold; Finished testing

Electrical (Pbar & EE Support)

  • Shunt Move
    • Progress
  • Shunt Repairs (6)
    • Removed and rebuilding
  • PS Maintenance
    • Started
  • Lens SCRs
  • Regulation Electronics
    • Started by EE
  • Ion Pump maintenance
    • Started; await fans
  • Deb Shunt re-arrangement
  • Deb Extraction kicker load adjustment

Other

  • Interlocks/Safety System (3)
  • Repair cathodic protection
  • Holes under buildings
    • Water dripping from tunnel ceiling
      • Contractors started
  • DCCT calibration
    • Started
  • Baking System checkout
  • RF & TWT filters maintenance
  • Lights Checkout
    • Fluorescents done
    • Emergency & Exit lights
  • Acc BPM repairs
    • Started
  • Activation Survey
  • Inductance checkout
    • Acc, Deb, Beamlines done
  • PM ODH Fans & Sumps
  • Tunnel Wireless
  • Install AP2 Leak Detector
  • SEM maintenance
  • Clogged sump line – Transport
    • Berm dug; Replaced entire line

Main Injector

  • 8 GeV line:
    • Replaced 3 vacuum valve solenoids
    • Installed multiwire
    • Leak checked
  • LEP Corrector coil replacement:
    • Complete
  • MI40 Lambertson beampipe replacement:
    • Setting up for replacement of both upstream and downstream beam pipes
  • Leak checking:
    • Complete at 527 and 614
  • Wet magnets from C0:
    • 5 magnets Shipped to IB2 for drying
  • MI10 kicker installation
    • Kicker installed
    • IPM needs repair
  • MI30 collimators
    • Making vacuum connections
    • Continuing buss and LCW moves
  • General Startup Plan
    • Turn on to pre-shutdown conditions.
      • Slip stacked beam to Pbar target
      • 2+5 mode for NuMI and Mixed mode
      • Use the old damper system
      • Use the pre-shutdown TLG modules
    • 2.2s cycle spacing
    • We would like to switch to multibatch slip stacking in 2 weeks
      • Need new TLG modules and commissioning activities must be complete
    • Measurements requiring dedicated time
      • MI8 line lattice measurements and matching (2 x 4 hrs.).
      • Aperture scans at MI30 (2 hrs. + problems)
      • Quad moves??? (8 hrs.)
      • Coalescing tune up (4 hrs.)
      • Slip Stacking tune up. (2 hrs.)
      • Tune up extraction positions and abort line (4 hrs.)
    • “Parasitic” Measurements
      • Commission new damper.
      • Collimator commissioning (MI8 and MI30).
      • Ring aperture scans with new software program.
      • General tune up
  • MI Status

Recycler

  • Finished or underway
    • New DCCT section installed, leak checked and under vacuum since Wednesday last week; pumping down
      • DCCT section has been aligned
      • This section will be baked
    • V303 moved
    • H308 moved
    • Need to check ALL LEP dipole positioning wrt. Beampipe
    • STC tank 7 leak check finished
      • Second attempt has been made; detected air but found no external leak
        • Sublimated TSPs at STCs 7,8,9,10.11,12
    • Trombone additional quadrupoles installation issues
      • HLRF cooperation
      • Installation has been completed
    • Added support to old DCCT
    • Bulk supply upgrade (on-going)
    • DCCT cable connected
    • Stochastic cooling: Photodiodes aligned
    • Leak check for the 800 transfer line: two leaks have been found so far
      • MW803 seal has been replaced; leak fixed!
        • MW803 survey requested and has been done
      • HP803 (large elliptical BPM) will be replaced with a standard elliptical unit
        • Leak check for replacement unit in progress
        • Survey of as-found has been done between gradient magnets
    • Measured rough positions of trombone elements with respect to surveyed elements
  • Rough Measurements of trim quads and LEP dipoles in the trombone section

  • Pelletron Status
    • Closed on Aug. 31 st
      • All maintenance finished except water
        • Replace filter and take water sample later on
    • Made access on Sept. 4 th to retrieve blue man-lift
    • U-bend mode running on Aug. 31 st
      • Very little gas activity (i.e. conditioning)
        • Although lost current was high at first
          • Appears to be back to nominal on Tuesday (Sept. 4)
    • Charging circuitry back to nominal (7.3 mA/kV)
    • Very good electric strength
      • Achieved 5 MV with only two vacuum bursts
    • Regulation loop parameters needed to be adjusted (new needles)
      • Will have to resave nominal file (for the full line) at start up
    • Got pulsing beam (full amplitude) to the collector with minimal tuning of a few correctors
    • Data taking with OTR started on August 31 st, in the afternoon
      • Camera timing was verified and the first pictures were acquired
    • Ran a good portion of the day on September 4 th
      • More to come
    • Issues:
      • We missed (during check out) what appears to be a broken resistor on the gun resistor chain
        • Only impact should be a lower resolution for BIASI
        • We won’t open the tank to fix it
          • Wait for a natural occurrence
      • One non-working BPM channel
        • Will be worked on today (September 5 th)
      • E50 is not working
    • Continue taking data with OTR
      • Intermittent data taking until we can go in the tunnel for YAG measurements
        • Primary goal (Arden): Polish data acquisition software and timing of the camera in preparation for YAG measurements
            • Secondary goal (Arden, Sasha, Lionel): Optics measurements, pepper-pot measurements (i.e. understand the drop of intensity seen on pepper-pot images)
    • Maybes (Lionel, Sasha): Check gun performance and calibration of the Gun CT
  • Recycler to do list
    • Magnet moves and survey /alignment pending calculations
    • Elevation survey after collimator installation
    • Electrical connection for new quadrupoles
    • Field check for new quadrupoles
    • LEP dipole field centering on beampipes
    • Vacuum maintenance
    • Kicker repair; timing verification
    • TOR330 repair
    • BPM calibration
    • Diagnostics maintenance
    • Laser link vacuum maintenance

Tevatron

  • Vacuum Issues
    • C4 very small leak some where can live with
    • D1 certified OK
    • E12-3 found leak had to pull magnet into isle to fix. Also possible bad spool at E12-1
    • A1 warmed up will look for leak after E1
    • D3 He to air leak
    • C0 upgrade issues with heater in MI magnets
  • Tevatron status as of September 14

SwitchYard

Continuing with maintenance work

NuMI

  • Complete HV101-1 magnet change out
  • Start installation of new Horn 2 RAW pressure gauges
  • Horn 2 RAW water leak repair
    • Develop repair procedures and ALARA plans
    • Working on upgrade to suction line ceramic
  • Develop plan to fill decay pipe with helium
    • Matt Cooper to lead engineering & procurement effort for helium filling system
    • ODH analysis for underground areas continues
    • Development of NuMI Note for changes Decay Pipe continues
    • Start installation of helium pipe to Absorber Hall starts Monday
  • Installed SWIC in muon alcove racks at Absorber Hall
  • Started PLC work on absorber skid
  • Target Pile Dehumidification Installation Project (40%)
    • Completed installation of chilled water piping & electrical conduit upper labyrinth
    • Completed installation of chilled water pipe & electrical conduit in T-Hall
    • Chilled water piping insulation completed in Target Hall
    • Completed electrical tie-ins to electrical panels for new installation during yesterdays power outage
    • Started construction dehumidifier deck east of work cell
    • Received chilled water strainer and condensate tank skid w/pumps

MiniBooNE

No work was completed this last week.

  • Reverse Horn Polarity (week of September 17)
    • Tale wipes
    • Reverse polarity of horn power supply
    • Test horn
  • Check Corrosion Tabs at 25m Absorber (27/28Aug07)
  • Change water DI Bottle (50%)
  • Dry out horn (09Aug07)
  • Check Leakage Current on Ion Pumps
  • Power Supply Maintenance
  • Toroid Work

MuCool Beamline Installation

  • Electrical Work (behind schedule)
    • Cable pull ongoing
      • All load cables have been pulled
        • EE support terminated the 4C#8 cable
      • Ongoing progress with instrumentation cables
        • ~ 9,000 ft of heliax cable pulled (90%)
    • Contractors will be working OT this weekend to finish up the cable pull
    • Next week:
      • Finish up contractor work (terminate 350/500MCM cables)
      • Instrumentation group, EE support, vacuum group will start terminating their cables
  • Mechanical Installation
    • N 2 line has been extended to fulfill MTA needs
  • Alignment
    • Alignment (behind schedule)
      • Network validation took longer than expected
        • Crew worked in Linac this week to collect more data
          • Network has been completed
    • EOI was sent to AMG regarding OT this weekend for MTA
      • Currently pending resolution
    • Next week: Rough alignment beam line until finished
      • Align the beam pipe sleeve
  • LCW work
    • Pipe fitters will be done today
    • Next week: Hydrostatic pressure test
      • Start making water connections to the magnets (?)
  • Vacuum work (behind schedule)
    • Waiting on alignment to proceed with spool piece connections
    • Next week: Installation of spool pieces as soon as possible
  • Proton Source start-up day: October 3 rd
    • MTA beam line installation has 10 days left!
    • There is a considerable amount of work remaining in order to reach the project goal
    • A reevaluation of the project goals is ongoing together with a revised schedule

Collider Detector at Fermilab

  • Silicon cooling leak repair
    • Both sides are leak tight enough to run
      • East leak rate reduced >1000
      • West a factor of a couple worse
    • Inspection of other cooling lines found no additional damage
      • All corrosion found was in the heat affected zones where cooling tubes are welded to manifold
      • Expect work to continue into next week
    • Close plug Thurs. or the following Mon. (9/24)
  • Detector survey 9/25
  • LBQ alignment 9/26-27
  • Other work ongoing

D-Zero

Work continues

Alignment

Completed Jobs

  • Recycler Task Manager: Oshinowo
    • 5527 Optically align Recycler DCCT after installation. Date: 8/13/2007 Complete
  • Tevatron Task Manager: Bocean
    • 5494 As found low beta quads and separators at B0 Date: 8/06/2007- 8/08/2007 Complete
    • 5495 As found low beta quads and separators at D0. Date: 8/09/2007- 8/14/2007 Complete
    • 5498 Measure rolls on 111 dipoles and quads in Tevatron (100 +) Complete
    • 5501 Un roll D16-1 quad when D-1 is warmed.(As Found) Date: 8/13/2007 Complete
    • 5591 As Founds of the Roman pots at D0 4 of 6 Date: 8/13/2007 85% Complete
    • 5598 A0 dump as found (no request for this job) Date: 8/23/2007 Complete
    • 5594 As found Schottkies, Separators and Beam pipe at A17 s Date: 8/24/2007 Complete
    • 5572 As found and re survey kicker replacement at F-17. Date: 8/28/2007 Complete
    • 5506 As found crystal collimator at E0. Date: 8/28/2007 Complete
    • 5506 level run through Tevatron tie rods. Date: 8/29/2007 Complete
    • As found TEL2 in transfer hall. Date: 9/18/2007 Complete
    • 5599 As found E 12-2 12-3 and 12-4 in preparation to pull E 12-3
  • Booster Task Manager: Oshinowo
    • 5534 Align 24 Booster Corrector Magnets.( Long 19 Long 17 ) Date: 8/13/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 21 Long 23 ) Date: 8/14/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long13 Long15 ) Date: 8/20/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 11 Long 9 ) Date: 8/21/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 7 Long 5 ) Date: 8/23/2007 Complete
    • 5534 Align 24 Booster Corrector Magnets ( Long 1 ) Date: 8/30/2007 Complete
  • Main Injector Task Manager: Oshinowo
    • 5563 AS found survey of Quadrupoles at 400, 402, 404 and the three Lambertson. Date: 8/13/2007- 8/15/2007 Complete
    • 5554 307, 308 Collimators: Base plate survey Date: 8/16/2007 Complete
    • 5554 307, 308 Collimators: Set travel and limit switches Date: 8/22/2007 Complete
    • 5596 Magnet 614 had to be moved please realign. Date: 8/23/2007 Complete
    • 5562 The beam pipe on the downstream side of the MI40 Lambertson. Date 8/29/2007 Complete
  • NuMI Task Manager: Bocean
    • 5602 As found the currently existing HV101-1 (spare EDBB017) dipole before replacement at MI-60. Date: 9/11/2007 Complete
  • CMS Task Manager: Oshinowo
    • 5516 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C Date: 8/09/2007 Complete
    • 5518 Survey the CMS F-pixel Half cylinder +Z2 in the Lab C Date: 9/06/2007 Complete
  • D0 Task Manager: Oshinowo
    • Stickmike survey of Muon system prior to opening of detector. This will include both CFs and EFs. Date: 8/06/2007 Complete
    • 5524 VStar or Laser Tracker survey of North/South EF pedestals with respect to the sidewalk monuments. Date: 8/07/2007 Complete
    • 5592 Elevation run in D0 before machine is opened. Date: 8/06/2007 Complete
  • P150 Task Manager: Oshinowo Late Request
    • 5593 We had to move several magnets and multiwires to make up the ends on the cans for vacuum connections. They are as follows: The multiwire cans at 708 and 710. B2 magnets at the interface of 710 - 711 and the interfaces of 709 - 710, the serial #'s are BDM 883-1 BDM 835-1 BDM 879-1 and BDM 397-1. Date: 9/07/2007 Complete
  • Linac Task Manager: Oshinowo
    • 5468 Stake out the MuCool beamline component locations in the Linac and MuCool Test Area (MTA). Date: 8/08/2007 Complete
  • P-Bar Task Manager: Oshinowo
    • 5488 As Found A4Q10 (LQA in Accumulator) has a water leak. Date: 8/09/2007 Complete
    • 5595 As Found An Accumulator multipole magnet has developed an leak. Date: 8/30/2007 Complete
    • 5488 Alignment A4Q10 (LQA in Accumulator) has a water leak. Date: 8/30/2007 Complete
    • 5595 Alignment An Accumulator multipole magnet has developed an leak. Date: 9/04/2007 Complete
    • 5540 The Debuncher DCCT (located between D1Q2 and D1Q3). Date: 9/12/2007 Complete
  • Nova Task Manager: Oshinowo
    • 5475- Stake out new Recycler injection line (MI10) Date: 8/27/2007 Complete
  • MTA Task Manager: Oshinowo
    • 5469 Vertical check of the MuCool beam line components. Date: 8/29/2007 Complete
    • Shoot MTA shielding blocks after re-stacking. Date: 9/04/2007 Complete

Next week’s schedule of work

  • Tevatron Task Manager: Bocean
    • 5499 Measure and re align 29 dipoles and quads that have new stands. Date: 9/04/2007- 9/07/2007
    • As found TEL1 in transfer hall. Date: 9/17/2007
    • 5603 As found and set new spool at E 12-1. Job must be done sometime this weekend
  • Elevation Run Floor points Manager: Kyle
    • 5528 Ring Wide elevation survey Date: 9/17/2007
  • Booster Task Manager: Oshinowo
    • 5534 Align 24 Booster Corrector Magnets Date: 9/04/2007
  • Main Injector Task Manager: Oshinowo
    • 5554 4 collimators in the Main Injector at the 301, 303, 307, and 308 locations. Date: 9/04/2007-9/07/2007
  • Nova Task Manager: Oshinowo
    • 5477-Stake out new Recycler RR30 quad positions for the Nova Project Date: 9/20/2007
    • 5476-Stake out new Recycler extraction line (MI30) for the Nova Project Accelerator upgrade Date: 9/20/2007
  • P-Bar Task Manager: Oshinowo
    • 5541 A new three-piece ceramic tube for the Debuncher Extraction Kicker (located between D6Q10 and D6Q9) will be installed. Want to assure the maximum alignment aperture is retained from production in Technical Division IB2 to installation on the stand (may have to take into account deflection of stand with/without kicker modules. Date: 9/14/2007
  • MTA Task Manager: Oshinowo
    • 5469 Rough alignment of the MuCool beam line components. Date: 9/15/2007 - 9/19/2007
    • 5538 Possibility to survey few 400MeV component(s) (2 quads?). There is a possibility that one has to remove temporary some 400MeV element(s) to make room for hauling the new Booster Corrector package. If that is the case, one needs to survey the components
  • CMS Task Manager: Oshinowo
  • CDF Task Manager: Greenwood
    • 5450 Measure elevation of CDF central detector Elevation Run. Date: 9/25/2007
  • NuMI Task Manager: Bocean
    • 5602 Align the HV101-1 magnet (the initial dipole EDBB020) after re-installation at MI-60. Date: 9/15/2007
  • ILC Task Manager: Bocean Not Scheduled

Alignment and Metrology Home page

Mechanical

Work in the tunnels continues.

Cryogenic System

Tevatron at LN2 temperature except A1 and E1 at room temperature

  • C4 and D1 were cooled down to LN2 temperature; no vacuum problems indicated
  • Pbar Debuncher Stochastic tanks and S/C Notch Filter Dewar are at LHe temperature
  • Tunnel magnet cryostat vacuums at A25 and D35 regions are of concern; D35 safety lead helium-atm leak was fixed by MSD/Cryo
  • 3 FRIG Mycom compressors (plus CDF & DZero) running with helium precooled to LN2 temperature flowing through magnet strings

Helium and Nitrogen Operations

  • CHL helium inventory management system operating on clean Helium Tank Farm, though A0 GHe Recovery being vented due to CHL Main Helium Purifier being offline while Screw compressors on Mobile Helium Purifier
  • CHL Nitrogen Reliquefier plant off and LN2 supply dependent on vendor deliveries to CHL LN2 Dewar tanks; deliveries are fine

FRIG Maintenance

  • 13 of 13 wet expander flywheels have been replaced.
  • 21 of 26 wet/dry expander overhauls of cold ends are completed
  • Various vacuum, instrumentation, and controls maintenance continuing.
  • Mycom compressor startup checkout in progress

CHL Maintenance

  • Coldbox-1:
    • Charcoal Adsorber west unit regeneration at high temperature almost finished
    • Oil Skid test run fine.
  • Nitrogen Reliquefier plant maintenance in progress:
    • R22 precooler skid compressor being inspected
    • TSI Compr 3000 hp motor replaced but bearing oil leak problems being addressed
  • Various compressors and facility maintenance in progress

Controls

This Week

  • Console Node OS Upgrades– Complete
  • States process port to Linux – Complete
  • Installation of 6 Booster VME Nodes with HRMs – Complete
  • NuMI AAT PLC Upgrade – In Progress
  • Linac Tank 1 Collimator Controls Upgrade – In Progress
  • MI-30 Collimator Controls Installation – In Progress
  • MADC II Updates continues
  • Routine maintenance continues

Next Week

  • 9/18/07
    • Database Disk Array Firmware Patch Installations
  • 9/19/07
    • Unix Webserver Maintenance
  • 9/20/07
    • Save/Restore and SDA Database Updates

Other Upcoming Disruptions

  • 9/27/07
    • Application and Device Database Updates

FESS

Sink Holes

  • BTE – SUMP DISCHARGE LINE REPAIR COMPLETED.
  • AP-0/AP40 – SUMP DISCHARGE LINE BREAK REPAIR COMPLETED


Casey’s Pond

  • NOW OPERATING On the 400HP PUMPS [flowing thru 14 inch strainer].
  • C-4 ICW Pumps not in operation unless required
  • ICW FLOW METER WIP. Expected completion/transfer back to the 16 inch strainer by/before 9/28/07

Poles

  • ALL of the19 poles are installed with conductors attached
  • Project completion is ahead of schedule
  • Power Up on 9/29/07 into MSS yard prior to site outage/MSS energizing on 10/3/07

TeV pond silt

  • ALL SECTORS COMPLETED FOR SILT TRANSFER.
  • SUCTION SCREEN EXTENSION FOR PO/INSTALLATION WIP [19 REQUIRED]. CRITICAL INSTALLATION AREAS TO BE COMPLETED BEFORE START UP

Other Jobs

  • CUB – 55LCW PUMP REPLACEMENTS WIP
  • TENTATIVE SITE ICW FLUSH SCHEDULED TO START SAT 9/29/07 THRU 10/06/07
  • ZEBRA MUSSEL TREATMENT SCHEDULED FOR MID-OCTOBER

Safety

  • Safety System Status
    • Upgrade M02 - M05 Enclosure Interlocks 65%
    • Upgrade M01 Enclosure Interlocks 10%
    • Upgrade Main Injector PLC Code and replace PLC Batteries 35%
    • Replace Safety System DEC talk with ATT Natural Voices 30%
    • Upgrade Meson Area Oxygen Monitors 0%
    • Replace Pbar Vault Key Tree Door Latches & add status LEDs 0%
    • Linux kernal 2.4 to 2.6 upgrade to Safety Data Acquisition System 100%
    • Route Key Tree release signals from MCR to Meson 100%
    • Install Radiation Monitor Cables for the Muon Test Area 100%
    • Replace oxygen monitor cells in the Tevatron 100%
  • A successful shutdown is getting our work done safely

Meson Oil Switch Replacement

  • MP Oil Switch
    • 9/24/07 @ 0700 – MW9 off 30 minutes to connect generator.
    • 9/25/07 @ 1600 – MW9 off 30 minutes to disconnect generator
    • 9/26/07 @ 0700 – MP9, MS7,PK120 off 30 minutes to connect generator
    • 9/26/07 – 10/5/07 – Running on Generator
    • 10/5/07 @ 1600 – MP9, MS7, PK120 off 30 minutes to disconnect generator
  • MXW Oil Switch
    • 10/8/07 @ 0700 – MAB & MS6 off 30 minutes to connect generator
    • 10/8/07 – 10/19/07 – Running on Generator
    • 10/19/07 @ 1600 – MAB & MS6 off 30 minutes to disconnect generator
  • MDE Oil Switch
    • 10/22/07 @ 0700 – Meson Detector Building (MDE) East off 30 minutes to connect generator. (MDW, MCC, LLRW have 5 minute outage)
    • 10/22/07 – 11/2/07 – Running on Generator
    • 11/02/07 @ 1600 – MDE off 30 minutes to disconnect generator. (MDW has a 5 minute outage)
    • 11/5/07 @ 0700 – Meson Cryo Central and Low Level Rad Waster (MCC&LLRW) have a 5 minute outage
    • 11/7/07 @ 0700 – LLRW has a 5 minute outage
    • 11/8/07 @ 1600 – LLRW has a 5 minute outage

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 09/14/2007    email Fermilab